# CRS-Diff: Controllable Generative Remote Sensing Foundation Model

### v2

@@ 4.4 Comparison and Analysis
Figure 4: Qualitative results of controlled images generated from a single condition. All of these conditions, with the exception of the Content condition, were accompanied by textual descriptions for participation in the bootstrapping process.
4.4.2 Compositional Image Generation.
In addition to the textual description, more bootstrap conditions were added to further guide the model towards generating more refined results. Single-conditional image generation. Figure 4 illustrates the image generation results achieved under the control of a single condition. To mitigate the risk of conflicts between the Content condition and the semantics of the textual description, resulting in inaccurate results, textual bootstrapping was added to all conditions except the Content condition. Among these, HED Boundary and Sketch contribute to intuitive image control by restricting the boundary and contour information of the generated image, making them suitable for mapping simulations; while Segmentation Mask, Road Map, and Content offer richer semantic information. Metadata control appears significantly more challenging, and efforts were made to select among the more prominent attributes, such as month and Cloud Cover. Under appropriate conditions, this is reflected. This information is effectively interpreted by CRS-Diff when combined with the appropriate prompts, influencing the generated results. Semantic segmentation and road detection are associated with two sub-tasks in the remote sensing image domain, thereby broadening the usability of CRS-Diff in the remote sensing domain. In contrast, Depthmap and MLSD provide limited information; Depthmap is capable of responding to the light and dark changes in remote sensing images under specific conditions, whereas MLSD conditions are more apt for architectural design simulation tasks and offer effective guidance for image synthesis.
Figure 5: More visual results of CRS-Diff under multiple condition control. The first three lines display the results of random combinations of some conditions and text descriptions, while the last two lines illustrate the results of the experiment after removing the text descriptions.
@@
Multi-conditional image generation. We conducted tests under multiple conditions, but it is worth noting that conflicts may arise among these conditions, and in the context of remote sensing images, even minor distortions can render the entire image meaningless, distinguishing from natural images. Therefore, in fig. 5, we present the results generated under the influence of multiple conditions, striving to minimize conflicts between these conditions. Utilizing textual guidance, CRS-Diff excels in generative capacity, controllability, and realism, successfully completing the synthesis of the target image. Despite the weak conditional information of our image (on the left side of the third row), a clear result is still achievable. However, without textual guidance, the quality of the synthesized image significantly degrades (the last two rows), evident in blurred lines and indistinct borders.
Table 2: Quantitative results comparisons among various CLIP models. The best results are in bold
Method IS $\uparrow$ FID $\downarrow$ CLIP Score ($\times$100) $\uparrow$ Zero-Shot classification OA $\uparrow$
CRS-Diff+CLIP(ViT-L) 14.96 108.01 19.89 53.40
CRS-Diff+CLIP(ViT-L*) 16.34 71.88 19.98 65.14
CRS-Diff+CLIP(ViT-bigG) 17.68 63.68 19.99 58.17
CRS-Diff+CLIP(ViT-bigG*) 18.39 50.72 20.33 67.38
Note: * Represents models fine-tuned on remote sensing data.
4.5 Ablation Analysis
For ablation experiments, our study discusses the Single image generation task mainly on the RSICD dataset, and we analyse the impact of CRS-Diff generation across various CLIP model bootstrapping approaches in table 2. The results show that both choosing a model with a larger number of parameters and fine-tuning on remote sensing images contribute significantly to the performance.
@@ 4.5 Ablation Analysis
Table 3: Experimental results comparison of road detection in the setting of different scaled datasets. Both the real and synthetic datasets consist of 5000 images. The best results are in bold
@@
Real 48.69 57.40 82.73 67.77
Real + Synthetic 48.98 58.89 81.75 67.92
4.6 Augmentation by our generated images
@@
For the condition generation phase, we posit that the generated image should have a sufficiently high correlation with the control labels as conditions. We aim for the generated image to encapsulate as much information from the conditioned image as possible and to offer data support for downstream tasks in related fields. Consequently, we incorporate experiments on the generated images to ascertain the pertinent performance of CRS-Diff. We consistently integrate synthetic dummy data into the training set of SGCN [51] for the road extraction task and assess it on the official test set. Experimental results in table 3 verify that the synthetic data can improve the performance of SGCN to some extent, which means that the generated RS images conditioned by road are beneficial to downstream road detection task.
5 Conclusion
@@
In this paper, we propose a new controllable RS image generation model (CRS-Diff), rooted in the foundational diffusion model architecture. Developed from the diffusion model framework, CRS-Diff enables high-quality remote sensing image generation. By integrating an optimized multi-conditional control mechanism, CRS-Diff can effectively synthesize multidimensional information, including text and images, guiding precise image generation and yielding highly accurate and controllable remote sensing images. This significantly broadens the control spectrum of the generation model, thus improving its adaptability to more complex application scenarios. Additionally, the generated images can help improve the performance of road detection task.
Supplementary Materials
Appendix 0.A Additional Training Details
We will detail the experimental process in separate sections, mainly including the two training phases, the datasets, and the corresponding processing utilized during training.
Single image generation. we utilized the corresponding weights from the Diffusers library and SD1.5, employing solely the RSICD dataset. Wherein during training, we randomly selected between 1 to 5 text descriptions for training data pairs. During the inference process, we generated 10 synthetic images for each text pair (excluding the untagged part of the original dataset), subsequently training a ResNet-18 model for evaluating the experimental results of IS, FID, according to the experimental criteria of Txt2Img. In zero-shot classification, we performed zero-sample classification on 2054 real images in the original test set and obtained the corresponding OA scores.
Compositional image generation. we initially reconstructed the relevant datasets (fMoW dataset and Million-AID dataset) to prepare for the subsequent experimental setup. Notably, in the caption field of the fMoW dataset, we added two easily comprehensible fields (Month and Cloud cover) to the text description and discarded them at a 20% probability. For the metadata, we encoded them into vectors of uniform length, utilizing the same embedding layer structure as the time step, and subsequently aggregated them to offer guidance. It is crucial to note that the impact of the control exerted by individual metadata varies significantly, with some fields, such as month, exhibiting a pronounced effect, while others display a weak influence, aligning with objective realities.
Appendix 0.B Additional Samples
0.B.1 Single Image Generation
Figure 6: Illustration of diversity results for single text generation of our method. Eight images are randomly generated corresponding to each given text description. No active effort is made to select visually superior synthetic images; the aforementioned images are all randomly generated from the model.
0.B.2 Compositional Image Generation
Figure 7: Illustration of diversity results under the control of a single condition of our method. Eight images are randomly generated corresponding to each given text description.
Figure 8: Illustration of diversity results for multiple conditions of our method. Eight images are randomly generated corresponding to each given text description.

### v3

@@ IV-D2 Single-condition image generation
Figure 7: Visual comparison results of generated RS images from single image condition. All the conditions are used together with the textual descriptions.
Quantitative analysis We compare CRS-Diff with ControlNet and Uni-ControlNet for quantitative evaluation on a test set of RSICD at a resolution of $512\times 512$. We randomly select one caption per image from the test set to be used as textual bootstrap information, obtaining 1k pieces of generated images for the quality evaluation. We employ the image decoupling method mentioned earlier to obtain more control conditions for constructing conditional data. Single condition generation (in addition to metadata) is restricted for quantitative evaluation. For Uni-ControlNet and ControlNet (Multi-ControlNet), the same dataset is used to train the conditional generation capabilities of the seven conditions. For HED, MLSD, Sketch, and Depthmap, we compute the SSIM of the generated images corresponding to the decoupled conditions. For the semantic segmentation mask and road map, we compute the mIoU. For the Content condition, considered as a text markup, we compute the CLIP Score using the CLIP model fine-tuned to the remotely sensed images. The specific results are shown in Tab. III. Our method achieves the best results on four metrics. Additionally, we calculate the FID metrics for the generated images, with specific results shown in Tab. IV. The experimental results demonstrate that CRS-Diff has excellent generative capabilities and quantitatively superior performance in most conditions compared to existing methods.
Figure 8: Visual comparison results of generated RS images under multiple condition control. Except for the content condition, all these conditions are used together with the textual descriptions.
IV-D3 Multi-condition image generation.
@@
We conducted tests under multiple conditions, but it is worth noting that conflicts may arise among these conditions, and in the context of RS images, even minor distortions can render the entire image meaningless, distinguishing from natural images. Therefore, in Fig. 8, we present the results generated under the influence of multiple conditions, striving to minimize conflicts between these conditions. Utilizing textual guidance, CRS-Diff excels in generative capacity, controllability, and realism, successfully completing the synthesis of the target image. Meanwhile, the generated images demonstrate sufficient diversity.
TABLE V: Ablations of Backbone model (ReB) and Feature Fusion network (FF).
ReB FF HED MLSD DepthMap Sketch Seg Road Content
(SSIM) (SSIM) (SSIM) (SSIM) (mIoU) (mIoU) (CLIP Score)
0.4212 0.6463 0.7219 0.4354 0.2748 0.3221 0.7784
0.4331 0.7834 0.6703 0.5594 0.2813 0.3741 0.8036
0.3845 0.7601 0.6575 0.4802 0.2452 0.3452 0.7966
0.4548 0.7455 0.8705 0.5676 0.2982 0.3787 0.8620
IV-E Ablation Analysis
For ablation experiments, we explore improvements in the structure of the multi-conditional control network and the method of injecting control information. We perform ablation experiments on CRS-Diff and its variants, analyzing the impact of replacing the backbone model (ReB) and the feature fusion approach (FF) on the generation quality and control effectiveness, respectively. We constructed a baseline based on the underlying Multi-ControlNet and executed the alteration approach sequentially, reporting the evaluation metrics of the different models, as shown in Tab. V. The pre-training of the backbone model and the incorporation of the feature fusion approach significantly improve the generation effect, enabling the model to generate remote sensing images with higher information density and realize the fusion of various control information types, further enhancing the model’s control generation capability.
@@ IV-E Ablation Analysis
TABLE VI: Experimental results comparison of road detection in the setting of different training datasets. Both the real and synthetic datasets consist of 1000 images.
@@
Real 52.84 59.26 85.51 69.49
Synthetic 50.67 57.07 84.24 67.54
Real + Synthetic 55.27 63.16 89.45 72.31
Figure 9: Visualisation comparison of SGCN model for road detection task under different settings of training datasets.
IV-F Application for downstream road detection task
@@
For the condition generation phase, we posit that the generated image should have a sufficiently high correlation with the control labels as conditions. We aim for the generated image to encapsulate as much information from the conditioned image as possible and to offer training data support for downstream tasks. Consequently, we incorporate experiments on the generated images to ascertain the pertinent performance of CRS-Diff. We consistently integrate synthetic dummy data into the training set of SGCN [50] for the road extraction task and assess it on the official test set. Tab. VI demonstrates the performance comparison of the SGCM method under different settings of training datasets. As can be seen, the synthetic training datasets can obtain almost the same performance as the real training dataset, which means that our CRS-Diff can simulate real images. By adding the synthetic dataset to the real dataset, the detection performance can be further significantly improved, indicating that the generated RS images conditioned by road are beneficial to the downstream road detection task. Additionally, Fig. 9 also visually compares the road extraction results under the three training conditions. The red boxes highlight areas where the differences are relatively large, revealing that the model trained with the synthetic dataset performs better in terms of continuity and completeness when dealing with a complex road network.
V Conclusion
@@
In this paper, we propose a new controllable generative RS foundation model (CRS-Diff). Developed from the diffusion model framework, CRS-Diff enables high-quality RS image generation. By integrating an optimized multi-conditional control mechanism, CRS-Diff can effectively synthesize multidimensional information, including text, metadata and images, guiding precise image generation and yielding highly accurate and controllable remote sensing images. This significantly broadens the control spectrum of the generation model, enhancing its adaptability to more complex application scenarios. Additionally, a comprehensive evaluation of existing multi-conditional generation models confirms CRS-Diff’s superior ability to generate remote sensing images under various conditions and its high controllability. This makes it suitable for a wide spectrum of use cases and enhances the performance of downstream tasks.